Bob finished his English assignment in 3/4 hours. Then he completed his chemistry assignment in 1/5 hours. What was the total amount of time Bob spent doing these two assignments?
1
Expert's answer
2020-09-14T19:16:54-0400

Bob finished his English assignment in "\frac{3}{4}" hours

Bob finished his chemistry assignment in "\frac{1}{5
}" hours


Total amount of time he spent=Time for English Assignment+Time for Chemistry Assignment

"=\dfrac{3}{4}+\dfrac{1}{5}\\
=\cfrac{3\times5+1\times4}{4\times 5}" ...............(On cross multiplication)

"=\cfrac{15+4}{20
}"

"=\cfrac{19}{20}hours"


Thus total time taken for Bob to complete these two assignments ="\frac{19}{20}" hours